OpenStudy (anonymous):

The total number of students and adults at a concert was 460. The cost of tickets was $12 per student and $16 per adult. The total amount of income from tickets was $6520. How many students and how many adults were at the concert?

OpenStudy (anonymous):

x+y=460 12x+16y=6520 where x is a student n y is an adult

OpenStudy (anonymous):

solving for x and y x=460-y 12(460-y)+16y=6520 5520-12y+16y=6520 4y=1000 y=250 x=210

OpenStudy (anonymous):

x+y=460 12x+16y=6520 y=460-x 12x+16(460-x)=6520 -4x+7360=6520 4x=840 x=210 y=250. There were 250 Adults.